OpenStudy (technodynamic):

Suppose f(x)=32+4x^2 Calculate the exact value(s) of x when f(x)=6.If there is more than one such value, enter all possible values as a comma separated list.

OpenStudy (anonymous):

6 = 32 + 4x^2 6-32 = 4x^2 -26 = 4x^2 -6.5 = x^2 you also get imaginary solutions ? cause no real solutions here..

OpenStudy (technodynamic):

How would I write that?

OpenStudy (anonymous):

imaginary : i = sqrt(-1) so : sqrt(-6.5) = x x = i * sqrt(6.5) x = +/- sqrt(6.5) * i

OpenStudy (technodynamic):

Okay thanks

OpenStudy (technodynamic):

I typed it in and it's not accepting it for some reason

OpenStudy (anonymous):

wait they are both good.. so maybe you dont have to use imaginary solutions and then there is no real solution ..